[12 pts] A supplier is willing to sell a container of orange juice for S()0.8q dollars. Meanwhile, consumers are willing to spend D(q) 7.8 0.5q dollars per container (this is demand). In this situation, q represents the number of containers of orange juice that are produced, in hundreds. A) How much is the producer willing to sell a container of orange juice for when they produce 600 containers? )comumer s wline eronaier ore iehow may o spend $3.50 for a c containers should be produced? C) Find the equilibrium quantity and price. Then find the revenue generated at equilibrium.
Answer #1

A) S(q) = 0.8q

q= 6 ( q is in hundreds, so for 600, q= 6)

S(q) = P = 0.8(6)= 4.8

Price that sellers are willing to sell at is 4.8

B) Now P= 3.5

3.5 = 0.8q

q= 4.375

Thus 437.5 should be produced

C) D(q)= P = 7.8 - 0.5q

At equilibrium D(q) = S(q)

7.8 - 0.5q = 0.8q

q= 6

P= 0.8(6)= 4.8

Thus equilibrium price is 4.8 and quantity is 600.

Revenue= Pq= 4.8(600)= 2880